OpenStudy (anonymous):

If α,β are the roots of \(x ^{2}-x+4=0\) , then \((3^{\alpha-1})(3^{\beta-1})=??\)

OpenStudy (anonymous):

@Hero @ikram002p @goformit100 would you kindly help me this question please?

OpenStudy (anonymous):

i guess it is \[\huge 3^{\alpha+\beta-2}\]

OpenStudy (anonymous):

the sum of the roots of a quadratic is \(-\frac{b}{a}\)

OpenStudy (anonymous):

sum of roots=1 product of roots=4 3^-1 = 1/3 ????

OpenStudy (anonymous):

which in your case is \[-\frac{b}{a}=-\frac{-1}{1}=1\]

OpenStudy (anonymous):

\[\huge 3^{\alpha+\beta-2}=3^{1-2}=\frac{1}{3}\] yes, what you got

OpenStudy (anonymous):

oh thank you so much. And I have another similar questions.